Graph y=f-1(x) if the graph of y=f(x)= 2x +5 restricted in the domain {x|-2 <_ x<_ 1.5|}
a. What is the range of the function?
b. What is the domain and range of its inverse

Answer by ikleyn(52958)

.
Graph y=f-1(x) if the graph of y=f(x)= 2x +5 restricted in the domain {x | -2 <= x <= 1.5|}
a. What is the range of the function?
b. What is the domain and range of its inverse
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

(a)  The range of this function is from its lover level  f(-2) = 2*(-2) + 5 = 1  

     to its upper level  f(1.5) = 2*1.5 + 5 = 8, inclusive.



(b)  The domain of its inverse is the same as the range of the direct function,

     which we determined in (a):  it is the interval  [1,8].



     The range of its inverse is the domain of the direct function, 
      which is given in the problem  [-2,1.5].

Solved: all questions are answered.